\begin{thebibliography}{10}

\bibitem{Ammari:2013:STN}
H.~Ammari, G.~Ciraolo, H.~Kang, H.~Lee, and G.~W. Milton.
\newblock Spectral theory of a {N}eumann--{P}oincar\'{e}--type operator and
  analysis of cloaking due to anomalous localized resonance.
\newblock {\em Arch. Ration. Mech. Anal.}, 208(2):667--692, 2013.

\bibitem{Ammari:2014:STN2}
H.~Ammari, G.~Ciraolo, H.~Kang, H.~Lee, and G.~W. Milton.
\newblock Spectral theory of a {N}eumann--{P}oincar\'{e}--type operator and
  analysis of cloaking by anomalous localized resonance {II}.
\newblock In {\em Inverse problems and applications}, volume 615 of {\em
  Contemp. Math.}, pages 1--14. Amer. Math. Soc., Providence, RI, 2014.

\bibitem{Ando:2019:SSN}
K.~Ando, Y.-G. Ji, H.~Kang, D.~Kawagoe, and Y.~Miyanishi.
\newblock {Spectral structure of the Neumann--Poincar\'e operator on tori}.
\newblock {\em Ann. Inst. Henri Poincar\'e C}, 36(7):1817--1828, 2019.

\bibitem{Ando:2016:PRF}
K.~Ando, H.~Kang, and H.~Liu.
\newblock Plasmon resonance with finite frequencies: a validation of the
  quasi-static approximation for diametrically small inclusions.
\newblock {\em SIAM J. Appl. Math.}, 76(2):731--749, 2016.

\bibitem{Basset:1893:TF}
A.~B. Basset.
\newblock On toroidal functions.
\newblock {\em American Journal of Mathematics}, 15(4):287--302, 1893.

\bibitem{Bonnetier:2012:PBG}
E.~Bonnetier and F.~Triki.
\newblock Pointwise bounds on the gradient and the spectrum of the
  {N}eumann-{P}oincar\'{e} operator: the case of 2 discs.
\newblock In {\em Multi-scale and high-contrast {PDE}: from modelling, to
  mathematical analysis, to inversion}, volume 577 of {\em Contemp. Math.},
  pages 81--91. Amer. Math. Soc., Providence, RI, 2012.

\bibitem{Bonnetier:2013:SPV}
E.~Bonnetier and F.~Triki.
\newblock On the spectrum of the {P}oincar\'{e} variational problem for two
  close-to-touching inclusions in 2{D}.
\newblock {\em Arch. Ration. Mech. Anal.}, 209(2):541--567, 2013.

\bibitem{Choi:2023:SAN}
D.~Choi, S.~P. Shipman, and M.~Lim.
\newblock {Spectral analysis of the Neumann-Poincar\'{e} operator for thin
  doubly connected domains}.
\newblock {\em arXiv:2309.02892}, 2023.

\bibitem{Chung:2014:CAL}
D.~Chung, H.~Kang, K.~Kim, and H.~Lee.
\newblock Cloaking due to anomalous localized resonance in plasmonic structures
  of confocal ellipses.
\newblock {\em SIAM J. Appl. Math.}, 74(5):1691--1707, 2014.

\bibitem{Cohl:2011:GHI}
H.~S. Cohl and D.~E. Dominici.
\newblock Generalized heine's identity for complex fourier series of binomials.
\newblock {\em Proceedings of the Royal Society A: Mathematical, Physical and
  Engineering Sciences}, 467(2126):333--345, 2011.

\bibitem{Grieser:2014:PEP}
D.~Grieser.
\newblock The plasmonic eigenvalue problem.
\newblock {\em Rev. Math. Phys.}, 26(3):1450005, 26, 2014.

\bibitem{Helsing:2017:CSN}
J.~Helsing, H.~Kang, and M.~Lim.
\newblock {Classification of spectra of the Neumann--Poincar\'e operator on
  planar domains with corners by resonance}.
\newblock {\em Ann. Inst. Henri Poincar\'e C}, 34(4):991--1011, 2017.

\bibitem{Jung:2018:SAN}
Y.~Jung and M.~Lim.
\newblock {Spectral analysis of the Neumann--Poincar\'{e} operator on the
  crescent-shaped domain and touching disks and analysis of plasmon resonance}.
\newblock {\em Nonlinear Analysis: Real World Applications}, 74:103951, 2023.

\bibitem{Kang:2017:SRN}
H.~Kang, M.~Lim, and S.~Yu.
\newblock {Spectral Resolution of the Neumann--Poincar\'e Operator on
  Intersecting Disks and Analysis of Plasmon Resonance}.
\newblock {\em Arch. Ration. Mech. Anal.}, 226(1):83--115, 2017.

\bibitem{Kang:2018:SPS}
H.~Kang and M.~Putinar.
\newblock Spectral permanence in a space with two norms.
\newblock {\em Rev. Mat. Iberoam.}, 34(2):621--635, 2018.

\bibitem{Kellogg:1929:FPT}
O.~D. Kellogg.
\newblock {\em {Foundations of Potential Theory}}, volume~31 of {\em Die
  Grundlehren der Mathematischen Wissenschaften}.
\newblock Springer-Verlag Berlin Heidelberg, 1929.

\bibitem{Khavinson:2007:PVP}
D.~Khavinson, M.~Putinar, and H.~S. Shapiro.
\newblock Poincar\'{e}'s variational problem in potential theory.
\newblock {\em Arch. Ration. Mech. Anal.}, 185(1):143--184, 2007.

\bibitem{Lebedev:1965:SFT}
N.~N. Lebedev.
\newblock {\em Special functions and their applications}.
\newblock Prentice-Hall, Inc., Englewood Cliffs, N.J., {English} edition, 1965.

\bibitem{Lim:2015:SBI}
M.~Lim.
\newblock Symmetry of a boundary integral operator and a characterization of a
  ball.
\newblock {\em Illinois J. Math.}, 45(2):537--543, 2001.

\bibitem{Mayergoyz:2005:ERN}
I.~D. Mayergoyz, D.~R. Fredkin, and Z.~Zhang.
\newblock Electrostatic (plasmon) resonances in nanoparticles.
\newblock {\em Phys. Rev. B}, 72:155412, Oct 2005.

\bibitem{Milton:2006:CEA}
G.~W. Milton and N.~A. Nicorovici.
\newblock On the cloaking effects associated with anomalous localized
  resonance.
\newblock {\em Proc. R. Soc. A}, 462(2074):3027--3059, 2006.

\bibitem{Neumann:1887:MAM}
C.~Neumann.
\newblock {\"{U}ber die Methode des Arithmetischen Mittels, Erste Abhandlung}.
\newblock {\em Abh. d. Kgl. S¨achs Ges. d. Wiss., IX and XIII}, 1887.

\bibitem{Ritter:1995:SEI}
S.~Ritter.
\newblock The spectrum of the electrostatic integral operator for an ellipsoid.
\newblock In {\em Inverse scattering and potential problems in mathematical
  physics ({O}berwolfach, 1993)}, volume~40 of {\em Methoden Verfahren Math.
  Phys.}, pages 157--167. Peter Lang, Frankfurt am Main, 1995.

\bibitem{Verchota:1984:LPR}
G.~Verchota.
\newblock Layer potentials and regularity for the {D}irichlet problem for
  {L}aplace's equation in {L}ipschitz domains.
\newblock {\em J. Funct. Anal.}, 59(3):572--611, 1984.

\bibitem{Zwillinger:2015:TIS}
D~Zwillinger, V.~Moll, I.S. Gradshteyn, and I.M. Ryzhik.
\newblock 8 - special functions.
\newblock In {\em Table of Integrals, Series, and Products (Eighth Edition)},
  pages 867--1013. Academic Press, Boston, 2015.

\end{thebibliography}

\documentclass[11pt,reqno]{article}
\usepackage[utf8]{inputenc}
\usepackage{booktabs} 
\usepackage{array} 
\usepackage{paralist} 
\usepackage{verbatim} 
\usepackage{amsthm}
\usepackage[table,xcdraw]{xcolor}
\usepackage[titletoc,toc,title]{appendix}
\usepackage{latexsym, amsmath, amssymb, a4, epsfig, color}
\usepackage{blindtext}
\usepackage{subcaption}
\usepackage[utf8]{inputenc}
\usepackage[export]{adjustbox}
\usepackage{wrapfig}
\usepackage{chngcntr}
\usepackage{apptools}
\usepackage{afterpage}
\usepackage{graphicx}
\usepackage{floatrow}

\usepackage{amssymb}
\usepackage{amsmath}
\usepackage[normalem]{ulem}

\usepackage{blkarray}
\usepackage{multirow}
\usepackage{mathtools}
\usepackage{graphicx}
\usepackage{arydshln}

\allowdisplaybreaks

\AtAppendix{\counterwithin{definition}{subsection}}
\AtAppendix{\counterwithin{theorem}{subsection}}

\graphicspath{}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{prop}[theorem]{Proposition}

\newcommand{\RR}{\mathbb{R}}
\newcommand{\CC}{\mathbb{C}}

\newcommand{\ZZ}{\mathbb{Z}}

\newcommand{\Om}{\Omega}
\newcommand{\ds}{\displaystyle}

\newcommand{\p}{\partial}
\newcommand{\pd}[2]{\frac {\p #1}{\p #2}}
\newcommand{\eqnref}[1]{(\ref {#1})}

\newcommand{\beq}{\begin{equation}}
\newcommand{\eeq}{\end{equation}}
\newcommand{\be}{\begin{equation*}}
\newcommand{\ee}{\end{equation*}}

\newcommand{\Kcal}{\mathcal{K}}
\newcommand{\Scal}{\mathcal{S}}

\newcommand{\la}{\langle}
\newcommand{\ra}{\rangle}

\numberwithin{equation}{section}
\numberwithin{figure}{section}

\begin{document}

\newcommand{\TheTitle}{Matrix representation of the Neumann--Poincar\'{e} operator for a torus}
\newcommand{\TheAuthors}{D. Choi and M. Lim}

\title{{\TheTitle}}

\author
{Doosung Choi\thanks{\footnotesize Department of Mathematics,  Louisiana State University, Baton Rouge, Louisiana 70803, USA ({dchoi@lsu.edu})}
}

\date{\today}
\maketitle

\begin{abstract}
We represent a matrix representation of the Neumann--Poincar\'{e} operator defined on the boundaries of a torus.  A torus is a doubly connected domain in three dimensions.  There is a well-known parametrization for the shape of the torus, the toroidal coordinate system.  Based on the coordinate system, we use toroidal harmonics to get an expansion of the NP operator for the torus.  Along with proper bases, the Neumann--Poincar\'{e} operator can be explicitly represented by an infinite matrix. 
 \end{abstract}
\noindent {\footnotesize {\bf Mathematics Subject Classification.} {
35J05; 	35P05; 
}}

\noindent {\footnotesize {\bf Keywords.} 
{Neumann--Poincar\'{e} operator; toroidal coordinate system; toroidal functions
}}

\section{Introduction}

Let $\Omega$ be a bounded and open region in $\RR^d$ $(d=2,3)$.  The Neumann--Poincar\'e (NP) operator associated with $\p \Omega$ is a boundary integral defined by
\beq\label{NP}
\Kcal_{\p\Om}^*[\varphi](x)= p.v.  \int_{\partial\Omega} \frac{\p \Gamma(x-y)}{\p \nu_x} \varphi(y)\, d\sigma(y),\quad x\in\p \Om,
\eeq
where $p.v.$ stands for the Cauchy principal value and $\nu_x$ is the outward unit normal vector at $x$.  Here, $\Gamma$ is the fundamental solution to the Laplacian in \eqnref{funsol}. The NP operator arises naturally when solving the Neumann or Dirichlet problems on interfaces.  Recently there is high interest in the spectral analysis of the NP operator due to its relation to nanophotonics and metamaterials: Plasmon resonance occurs near the eigenvalues of the NP operator and cloaking due to anomalous localized resonance (CALR) occurs at the accumulation point of eigenvalues, respectively \cite{Ammari:2013:STN, Ando:2016:PRF,Bonnetier:2012:PBG,Bonnetier:2013:SPV, Grieser:2014:PEP, Mayergoyz:2005:ERN, Milton:2006:CEA}. 

Although the NP operator is not symmetric on $L^2(\p\Omega)$ unless $\Omega$ is a disk or a ball \cite{Lim:2015:SBI}, it can be symmetrized on the Sobolev space $H^{-1/2}_0(\p\Omega)$ using the Plemelj's symmetrization principle \cite{Khavinson:2007:PVP}, where $H^{-1/2}_0(\p\Omega)$ is a space $H^{-1/2}(\p\Omega)$ with the mean zero condition.  The spectrum of the NP operator, denoted by $\sigma(\Kcal_{\p\Om}^*)$,  consists of absolutely continuous spectrum, singularly continuous spectrum, and pure point spectrum. More specifically,  $\sigma(\Kcal_{\p\Om}^*)\setminus \{1/2\}$ is contained in $(-1/2,1/2)$ \cite{Kang:2018:SPS,Kellogg:1929:FPT, Verchota:1984:LPR}. 
If $\Omega$ has $C^{1,\alpha}$ boundary with some $\alpha\in(0,1)$, then the NP operator is compact and thus $\sigma(\Kcal_{\p\Om}^*)$ is a sequence that accumulates to $0$.  One can easily prove that $\{0, 1/2\}$ is the spectrum of the NP operator if $\Omega$ is a disk.  Complete sets of $\sigma(\Kcal_{\p\Om}^*)$ for a sphere, an ellipse \cite{Neumann:1887:MAM},  and an ellipsoid are already known \cite{Ritter:1995:SEI}.  In two dimensions,  the twin spectrum relation for the NP operator holds \cite{Mayergoyz:2005:ERN}. The explicit spectrum of the NP operator has been revealed for various shapes.  For example,  Kang et al.  investigated a spectral resolution of the NP operator on intersecting disks with corners using the bipolar coordinate system \cite{Kang:2017:SRN}.  Jung and Lim studied for touching disks and crescent domains \cite{Jung:2018:SAN} and Helsing et al.  classified the spectra of the NP operator on a planar domain with corners \cite{Helsing:2017:CSN}. 

For doubly connected domains,  Ammari et al.  developed spectrum of concentric disks and spheres, and derived a necessary and sufficient condition for CALR to take place \cite{Ammari:2013:STN, Ammari:2014:STN2}, After that, Chung et al.  computed eigenvalues and eigenfunctions of the NP operator associated with confocal ellipses \cite{Chung:2014:CAL}.  Recently, Choi et al. showed that spectrum of the NP operator for an arbitrary-shaped, planar, and doubly connected domain approaches to $[-1/2,1/2]$ as thickness of the shape tends to zero \cite{Choi:2023:SAN}.  In addition, Ando et el.  proved that the NP operator defined on the boundary of a torus has infinitely many positive and negative eigenvalues \cite{Ando:2019:SSN}. To establish this, the NP operator is decomposed into self-adjoint, compact operators, showing that both positive and negative eigenvalues exist in their numerical range.  This is the first result that identified the torus as the three dimensional surface with infinitely many negative eigenvalues.



We explicitly find the expansion the NP operator of a torus by using the toroidal harmonics.  We investigate spectral analysis of those doubly connected domains in three-dimensions.  We acquire harmonics expansions of the NP operators from the suggested coordinate systems.

In this paper, we explore the basis for the density function of the NP operator using the toroidal coordinate system as introduced in \cite{Ando:2019:SSN}. Furthermore, we employ the toroidal functions, which are a specialized form of the associated Legendre functions. We then decompose the NP operator analytically in the form of infinite matrices.

\section{Preliminary}

\subsection{Layer Potentials}

The fundamental solution $\Gamma$ to the Laplacian has the form
\beq\label{funsol}
\Gamma(x) = 
\begin{dcases}
\frac{\ln|x|}{2\pi},  \quad  d=2,\\[1mm]
\frac{1}{4\pi |x|}, \quad  d=3.
\end{dcases}
\eeq
For a density function $\varphi\in L^2(\p\Om)$, we define the single layer potential associated with $\p\Om$ as
\begin{align*}
\ds&\Scal_{\p\Om}[\varphi](x)= \int_{\p \Om} \Gamma(x-y) \, \varphi(y)\, d\sigma(y),\quad x\in\RR^d
\end{align*}
and the single layer potential satisfies the following jump relations on the boundary \cite{Verchota:1984:LPR}:
\begin{align}\label{S_nd}
\ds\frac{\partial}{\partial\nu}\Scal_{\p\Om}[\varphi]\Big|_{\pm}&=\left(\pm\frac{1}{2}I+\Kcal_{\p\Om}^*\right)[\varphi]\quad\text{on }\partial\Omega
\end{align}
with the NP operator $\Kcal_{\p\Om}^*$ defined in \eqnref{NP}, and we have from \eqnref{funsol},
$$
\Kcal_{\p\Om}^*[\varphi](x)= p.v. \frac{1}{\omega_d} \int_{\partial\Omega} \frac{\left\la x-y,\nu_x\right\ra}{|x-y|^d}\, \varphi(y)\, d\sigma(y),\quad x\in\p \Om,
$$
where $\omega_d = 2\pi$ if $d=2$ and $\omega_d = 4\pi$ if $d=3$.

\subsection{Toroidal coordinate system}

Toroidal coordinates are a three-dimensional orthogonal system that is formed by rotating the two-dimensional bipolar coordinate system around the axis that divides its two focal points. The two foci in bipolar coordinates lies on the unit circle in the $xy$-plane of the toroidal system, where $z$-axis is the axis of the rotation. The unit circle is called the focal ring or the reference circle.

We define $(x,y,z)$ as the Cartesian coordinate system. The toroidal coordinate system $(\tau,\phi,\sigma)$ relates with the Cartesian coordinate system as follows:
\begin{align}\label{toroidal}
x = \frac{a\sinh\tau \cos\phi}{\cosh\tau - \cos\sigma}, \quad y = \frac{a\sinh\tau \sin\phi}{\cosh\tau - \cos\sigma}, \quad z = \frac{a\sin\sigma}{\cosh\tau - \cos\sigma},
\end{align}
where the coordinate ranges are $\tau>0$, $0\le \phi < 2\pi$, and $0\le \sigma < 2\pi$.  The coordinate surface $\tau = \tau_0$ with a constant $\tau_0>0$ represents a torus. 

It is well-known that the normal derivative and the tangential derivatives of a function with toroidal coordinates satisfy
\begin{align}
&\frac{\p u}{\p \nu} = \frac{\cosh\tau - \cos\sigma}{a} \frac{\p u}{\p \tau}, \label{nd}\\[2mm]
&\frac{\p u}{\p T_\sigma} = \frac{\cosh\tau - \cos\sigma}{a} \frac{\p u}{\p \sigma}, \notag\\[2mm]
&\frac{\p u}{\p T_\phi} = \frac{\cosh\tau - \cos\sigma}{a\sinh\tau} \frac{\p u}{\p \phi}. \notag
\end{align}

\subsection{Toroidal harmonics}

One advantage of the toroidal coordinates is that we can apply the separation of variable method to solve the Laplace equation. The solution of the Laplace equation $\Delta u = 0$ has the form
$$
u(\tau,\phi,\sigma) = (\cosh\tau - \cos\sigma)^{\frac{1}{2}} \, u_1(\tau) \, u_2(\phi) \, u_3(\psi).
$$
In fact, the toroidal harmonics of the above form are
\beq\label{GH}
\begin{aligned}
&G_{mn}(\tau,\phi,\sigma) = (\cosh\tau - \cos\sigma)^{\frac{1}{2}} \, Q_{n-1/2}^m (\cosh\tau) \, e^{im\phi} \, e^{in\sigma},\\
&H_{mn}(\tau,\phi,\sigma) = (\cosh\tau - \cos\sigma)^{\frac{1}{2}} \, P_{n-1/2}^m (\cosh\tau) \, e^{im\phi} \, e^{in\sigma},
\end{aligned}
\eeq
where $P_\nu^\mu$ and $Q_\nu^\mu$ $(\nu,\mu\in\CC)$ are the associated Legendre function of the first and second kind, respectively.  In particular, if $\nu=n-1/2$ and $\mu=m$ for $m,n\in\ZZ$, we call $P_{n-1/2}^m$ and $Q_{n-1/2}^m$ the toroidal or ring functions.  The following formulas are the integral representation of the toroidal functions \cite{Zwillinger:2015:TIS}:
\begin{align*}
&P_{n-1/2}^m(\cosh\tau) = \frac{(-1)^m}{2\pi} \frac{\Gamma(n+1/2)}{\Gamma(n-m+1/2)} \int_0^{2\pi} \frac{\cos m\theta}{(\cosh\tau + \sinh\tau \cos \theta)^{n+\frac{1}{2}}} d\theta,\\
&Q_{n-1/2}^m(\cosh\tau) = (-1)^m \frac{\Gamma(n+1/2)}{\Gamma(n-m+1/2)} \int_0^\infty \frac{\cosh mt}{(\cosh\tau + \sinh\tau \cosh t)^{n+\frac{1}{2}}} dt.
\end{align*}

$G_{mn}$ is harmonic in $\RR^3$ except for the $z$-axis and called the internal toroidal harmonics.  $H_{mn}$ is harmonic in $\RR^3$ except for the reference circle $z=0$, $x^2+y^2=1$ and called the external toroidal harmonics.  We refer \cite{Basset:1893:TF, Zwillinger:2015:TIS} for further details on the toroidal harmonics and the toroidal functions.

\section{Matrix representation of the NP operator for a torus}

Let us consider a three dimensional region, i.e., $d=3$.  We define a function 
$$
u(x) = \Scal_{\p\Om}[\varphi](x)
$$ 
for $\varphi \in H_0^{-\frac{1}{2}}(\p \Omega)$.  Then $u$ is the solution of the system
\beq\label{problem}
\begin{cases}
\ds \Delta u = 0 \qquad&\mbox{in }\mathbb{\RR}^3 \setminus \p \Omega, \\[2mm]
\ds u\big|_+ = u\big|_- \qquad&\mbox{on }\p \Om,\\[2mm]
\ds \pd{u}{\nu}\Big|_+ - \pd{u}{\nu}\Big|_- = \varphi \qquad&\mbox{on }\p \Om,\\[2mm]
\ds u(x) = O(|x|^{-2}) \qquad&\mbox{as }|x|\to\infty.
\end{cases}
\eeq
Let $\Omega$ be a torus with the boundary satisfying $\{ (\tau,\phi,\sigma)\in\RR^3:\tau=\tau_0 \}$. According to the definition of the toroidal harmonics in \eqnref{GH},  the following $u_{mn}$ is a solution of the system \eqnref{problem}:
\beq\label{solution}
u_{mn}(\tau,\phi,\sigma)=
\begin{cases}
\ds P_{n-1/2}^m(\cosh\tau_0) \, G_{mn}(\tau,\phi,\sigma) \qquad&\mbox{in } \Omega, \\[4mm]
\ds Q_{n-1/2}^m(\cosh\tau_0) \, H_{mn}(\tau,\phi,\sigma) \qquad&\mbox{in }\mathbb{\RR}^3 \setminus \overline{\Omega}.
\end{cases}
\eeq
Here, $\Omega = \{ (\tau,\phi,\sigma)\in\RR^3:\tau>\tau_0 \}$ and $\mathbb{\RR}^3 \setminus \overline{\Omega} = \{ (\tau,\phi,\sigma)\in\RR^3:\tau<\tau_0 \}$. There exist a density function $\varphi_n^m = \varphi_n^m(\phi,\sigma) \in H_0^{-\frac{1}{2}}(\p \Omega)$ such that
$$
\varphi_n^m = \pd{u_{mn}}{\nu}\Big|_{\tau=\tau_0^-} - \pd{u_{mn}}{\nu}\Big|_{\tau=\tau_0^+}
\quad 
\mbox{and}
\quad 
u_{mn}(x) = \Scal_{\p\Om}[\varphi_n^m](x).
$$
Using \eqnref{nd}, we obtain
\begin{align}
\varphi_n^m 
&= \pd{u_{mn}}{\nu}\Big|_{\tau=\tau_0^-} - \pd{u_{mn}}{\nu}\Big|_{\tau=\tau_0^+} \notag\\
&= \frac{\cosh\tau_0 - \cos\sigma}{a} \Bigg[Q_{n-1/2}^m(\cosh\tau_0) \pd{ H_{mn}(\tau,\phi,\sigma)}{\tau}\Big|_{\tau=\tau_0} - P_{n-1/2}^m(\cosh\tau_0) \pd{G_{mn}(\tau,\phi,\sigma)}{\tau}\Big|_{\tau=\tau_0} \Bigg].\label{basis}
\end{align}
As $H_{mn}$ and $G_{mn}$ have a relation
$$
H_{mn}(\tau,\phi,\sigma) = G_{mn}(\tau,\phi,\sigma) \frac{P_{n-1/2}^m(\cosh\tau)}{Q_{n-1/2}^m(\cosh\tau)},
$$
the derivative of the product follows that
\beq\label{Hderiv}
\pd{ H_{mn}(\tau,\phi,\sigma)}{\tau} = \frac{\p G_{mn}(\tau,\phi,\sigma)}{\p \tau} \frac{P_{n-1/2}^m(\cosh\tau)}{Q_{n-1/2}^m(\cosh\tau)}  +  G_{mn}(\tau,\phi,\sigma)  \frac{d}{d \tau}\Bigg[ \frac{P_{n-1/2}^m(\cosh\tau)}{Q_{n-1/2}^m(\cosh\tau)} \Bigg].
\eeq
Let $\eta = \cosh\tau$.  Then the toroidal functions have the Wronskian relation as the next lemma.
\begin{lemma}[\cite{Lebedev:1965:SFT}]\label{Wron}
For $z = \cosh\tau$, the toroidal functions satisfy the Wronskian relation
\be
Q_{n-1/2}^m(z)\frac{d}{d z} P_{n-1/2}^m(z) - P_{n-1/2}^m(z)\frac{d}{d z}Q_{n-1/2}^m(z)
=
\frac{(-1)^m}{z^2-1} \frac{\Gamma(n+m+1/2)}{\Gamma(n-m+1/2)}
\ee
for all $m,n\in\ZZ$.
\end{lemma}
By elliminating common terms in \eqnref{basis} from \eqnref{Hderiv},  the density function holds that
\begin{align*}
\varphi_n^m(\phi,\sigma)
&= \frac{\cosh\tau_0 - \cos\sigma}{a} Q_{n-1/2}^m(\cosh\tau_0) \, G_{mn}(\tau_0,\phi,\sigma)  \frac{d}{d \tau}\Bigg[ \frac{P_{n-1/2}^m(\cosh\tau)}{Q_{n-1/2}^m(\cosh\tau)} \Bigg]\Bigg|_{\tau=\tau_0}\\
&= \frac{\cosh\tau_0 - \cos\sigma}{a} Q_{n-1/2}^m(\cosh\tau_0) \, G_{mn}(\tau_0,\phi,\sigma)  \frac{d \cosh\tau}{d \tau}\Bigg|_{\tau=\tau_0} \frac{d}{d z} \Bigg[ \frac{P_{n-1/2}^m(z)}{Q_{n-1/2}^m(z)} \Bigg]\Bigg|_{z=\cosh(\tau_0)},\\
&= \frac{\cosh\tau_0 - \cos\sigma}{a} Q_{n-1/2}^m(\cosh\tau_0) \, G_{mn}(\tau_0,\phi,\sigma) \sinh\tau_0 \\
& \quad \times \frac{Q_{n-1/2}^m(\cosh(\tau_0))\frac{d}{d z} P_{n-1/2}^m(z)  |_{z=\cosh(\tau_0)}- P_{n-1/2}^m(\cosh(\tau_0))\frac{d}{d z}Q_{n-1/2}^m(z) |_{z=\cosh(\tau_0)} }{\Big[Q_{n-1/2}^m(\cosh(\tau_0))\Big]^2},\\
&= \frac{(-1)^m}{a \sinh\tau_0} \frac{\Gamma(n+m+1/2)}{\Gamma(n-m+1/2)} (\cosh\tau_0 - \cos\sigma)^{\frac{3}{2}} e^{im\phi} \, e^{in\sigma},
\end{align*}
where the last formula holds from \eqnref{GH} and Lemma \ref{Wron}. Thus, we get the following proposition.

\begin{prop}\label{density_formula}
The density function $\varphi_n^m$ for the solution \eqnref{solution} of the system \eqnref{problem} has the form
\begin{align*}
\varphi_n^m(\phi,\sigma)
&= \Gamma_{mn} (\cosh\tau_0 - \cos\sigma)^{\frac{3}{2}} \, e^{im\phi} \, e^{in\sigma} \quad \mbox{for } m,n\in\ZZ,
\end{align*}
where $\Gamma_{mn}$ is the constant,
\beq\label{constantC}
\Gamma_{mn} = \frac{(-1)^m}{a \sinh\tau_0} \frac{\Gamma(n+m+1/2)}{\Gamma(n-m+1/2)}.
\eeq
\end{prop}

Let us find the spectral decomposition of the NP operator.  As $\Omega = \{ (\tau,\phi,\sigma)\in\RR^3:\tau>\tau_0 \}$, the jump formula \eqnref{S_nd} gives
$$
\frac{\partial}{\partial\nu}\Scal_{\p\Om}[\varphi_n^m](\tau,\phi,\sigma)\Big|_{\tau=\tau_0^\mp} = \pm\frac{1}{2} \varphi_n^m(\phi,\sigma) + \Kcal_{\p\Om}^*[\varphi_n^m](\phi,\sigma) \quad \mbox{for } m,n\in\ZZ.
$$
Then \eqnref{nd}, \eqnref{solution},  and Proposition \ref{density_formula} imply that
\begin{align*}
&\Kcal_{\p\Om}^*[\varphi_n^m](\phi,\sigma)\\
&= \frac{1}{2} \left[ \frac{\partial}{\partial\nu}\Scal_{\p\Om}[\varphi_n^m](\tau, \phi,\sigma)\Big|_{\tau=\tau_0^-} + \frac{\partial}{\partial\nu}\Scal_{\p\Om}[\varphi_n^m](\tau, \phi,\sigma)\Big|_{\tau=\tau_0^+} \right]\\
&= \frac{1}{2} \left[ Q_{n-1/2}^m(\cosh\tau_0) \frac{\partial}{\partial\nu} H_{mn}(\tau,\phi,\sigma)\Big|_{\tau=\tau_0} + P_{n-1/2}^m(\cosh\tau_0) \frac{\partial}{\partial\nu} G_{mn}(\tau,\phi,\sigma)\Big|_{\tau=\tau_0} \right]\\
&= \frac{\cosh\tau_0 - \cos\sigma}{2a} \left[ Q_{n-1/2}^m(\cosh\tau_0) \frac{\partial}{\partial\tau} H_{mn}(\tau,\phi,\sigma)\Big|_{\tau=\tau_0} + P_{n-1/2}^m(\cosh\tau_0) \frac{\partial}{\partial\tau} G_{mn}(\tau,\phi,\sigma)\Big|_{\tau=\tau_0} \right]\\
&= \frac{\cosh\tau_0 - \cos\sigma}{2a} \Bigg[ Q_{n-1/2}^m(\cosh\tau_0) \frac{\partial}{\partial\tau} \left( (\cosh\tau - \cos\sigma)^{\frac{1}{2}} \, P_{n-1/2}^m (\cosh\tau) \right)\Big|_{\tau=\tau_0} \\
&\qquad\qquad\qquad\qquad + P_{n-1/2}^m(\cosh\tau_0) \frac{\partial}{\partial\tau} \left( (\cosh\tau - \cos\sigma)^{\frac{1}{2}} \, Q_{n-1/2}^m (\cosh\tau) \right)\Big|_{\tau=\tau_0} \Bigg] \, e^{im\phi} \, e^{in\sigma}\\
&= \frac{(z_0 - \cos\sigma) \sinh\tau_0}{2a} \Bigg[ Q_{n-1/2}^m(z_0) \left( \frac{1}{2} (z_0 - \cos\sigma)^{-\frac{1}{2}} \, P_{n-1/2}^m (z_0) + (z_0 - \cos\sigma)^{\frac{1}{2}} \, \frac{\partial}{\partial z} P_{n-1/2}^m (z_0)  \right) \\
&\qquad\qquad\qquad\qquad\qquad + P_{n-1/2}^m(z_0) \left( \frac{1}{2} (z_0 - \cos\sigma)^{-\frac{1}{2}} \, Q_{n-1/2}^m (z_0) + (z_0 - \cos\sigma)^{\frac{1}{2}} \, \frac{\partial}{\partial z}  Q_{n-1/2}^m (z_0)  \right) \Bigg] \, e^{im\phi} \, e^{in\sigma},\\
&= \frac{(z_0 - \cos\sigma)^{\frac{3}{2}} \sinh\tau_0}{2a} \Bigg[ (z_0 - \cos\sigma)^{-1} P_{n-1/2}^m(z_0) \, Q_{n-1/2}^m (z_0) \\
&\qquad\qquad\qquad\qquad\qquad + Q_{n-1/2}^m(z_0)   \, \frac{\partial}{\partial z} P_{n-1/2}^m (z_0) + P_{n-1/2}^m(z_0) \frac{\partial}{\partial z}  Q_{n-1/2}^m (z_0) \Bigg] \, e^{im\phi} \, e^{in\sigma},
\end{align*}
where the last equality holds from $\frac{\partial}{\partial\tau} = \frac{d z}{d\tau} \frac{\partial}{\partial z}$ with $z = \cosh\tau$ and $z_0 = \cosh\tau_0$.  Hence, we have
\begin{align}\label{K1}
\Kcal_{\p\Om}^*[\varphi_n^m](\phi,\sigma)
&= \frac{(-1)^m \sinh^2\tau_0}{2} \frac{\Gamma(n-m+1/2)}{\Gamma(n+m+1/2)} \, \varphi_n^m(\phi,\sigma) \notag\\
&\quad \times\Bigg[ \frac{P_{n-1/2}^m (z_0) \, Q_{n-1/2}^m (z_0)}{\cosh\tau_0 - \cos\sigma} + Q_{n-1/2}^m(z_0)  \, \frac{\partial}{\partial z} P_{n-1/2}^m (z_0) + P_{n-1/2}^m (z_0) \, \frac{\partial}{\partial z}  Q_{n-1/2}^m (z_0) \Bigg].
\end{align}
We need the following lemma of which gives the Fourier expansion of a function in \eqnref{K1}.

\begin{lemma}[\cite{Cohl:2011:GHI}]\label{GHI}
For $\tau_0>0$ and $0\le \sigma < 2\pi$,  we have
\beq
\frac{\sinh\tau_0}{\cosh\tau_0- \cos \sigma}
= 1 + 2 \sum_{k=1}^\infty e^{-k\tau_0} \cos(k\sigma)
= \sum_{k=-\infty}^\infty e^{-|k|\tau_0} e^{ik\sigma}.
\eeq
\end{lemma}

\begin{lemma}\label{toroidrelation}
For $m,n\in\ZZ$, the toroidal function satisfies the relations
$$
P^{-m}_{n-\frac{1}{2}}(z) = \frac{\Gamma(n-m+1/2)}{\Gamma(n+m+1/2)} P_{n-1/2}^m(z).
$$
\end{lemma}

\begin{lemma}\label{deriv}
For $m,n\in\ZZ$ and $z>1$, the toroidal functions satisfy the derivative formulas
\begin{align*}
& P_{n-1/2}^{-m}(z) = (n+1/2+m)P_{n+1/2}^{-m}(z) - (n+1/2)  z_0 P_{n-1/2}^{-m}(z),\\
& Q_{n-1/2}^m(z) = (n+1/2-m) Q_{n+1/2}^m(z) - (n+1/2)  z_0 Q_{n-1/2}^m(z).
\end{align*}
\end{lemma}


\begin{prop}\label{Kexpansion}
Let $\{\varphi_n^m\}_{m,n\in\ZZ}$ be the basis for the density functions.
The NP operator for tori has the expansion,
$$
\Kcal_{\p\Om}^*[\varphi_n^m]= D_{nn}^{(m)} \varphi_n^m + \sum_{l=-\infty}^\infty R_{nl}^{(m)} \varphi_l^m,
$$
where $D^{(m)} = \left[ D_{nl}^{(m)} \right]_{n,l\in\ZZ}$ is a diagonal matrix and $R^{(m)} = \left[ R_{nl}^{(m)} \right]_{n,l\in\ZZ}$ defined by
\begin{align*}
D_{nn}^{(m)}
&= \frac{(-1)^m}{2} \Bigg[ (n-m+1/2) Q_{n+1/2}^m(z_0) P_{n-1/2}^{-m}(z_0) + (n+m+1/2) Q_{n-1/2}^m(z_0) P_{n+1/2}^{-m}(z_0) \notag\\
&\qquad\qquad\quad - (2n+1)  z_0 Q_{n-1/2}^m(z_0) P_{n-1/2}^{-m}(z_0) \Bigg],\\
R_{nl}^{(m)}
&= \frac{(-1)^m \sinh\tau_0}{2e^{|l-n|\tau_0}} Q_{n-1/2}^m (z_0) P_{l-\frac{1}{2}}^{-m}(z_0).
\end{align*}
\end{prop}
\begin{proof}
Let us define
\begin{align*}
D_{nn}^{(m)} 
&= \frac{(-1)^m \sinh^2\tau_0}{2} \frac{\Gamma(n-m+1/2)}{\Gamma(n+m+1/2)} \, \Bigg[ Q_{n-1/2}^m(z_0)  \, \frac{\partial}{\partial z} P_{n-1/2}^m (z_0) + P_{n-1/2}^m (z_0) \, \frac{\partial}{\partial z}  Q_{n-1/2}^m (z_0) \Bigg].
\end{align*}
Applying Lemma \ref{GHI} to \eqnref{K1} generates the spectral decomposition of the NP operator:
\begin{align*}
&\Kcal_{\p\Om}^*[\varphi_n^m](\phi,\sigma)\\
&= D_{nn}^{(m)} \varphi_n^m(\phi,\sigma) +  \frac{(-1)^m \sinh\tau_0}{2} \frac{\Gamma(n-m+1/2)}{\Gamma(n+m+1/2)} \, P_{n-1/2}^m (z_0) \, Q_{n-1/2}^m (z_0) \, \frac{\sinh\tau_0}{\cosh\tau_0 - \cos\sigma} \, \varphi_n^m(\phi,\sigma) \\
&= D_{nn}^{(m)} \varphi_n^m(\phi,\sigma) +  \frac{(-1)^m \sinh\tau_0}{2} \frac{\Gamma(n-m+1/2)}{\Gamma(n+m+1/2)} \, P_{n-1/2}^m (z_0) \, Q_{n-1/2}^m (z_0) \, \sum_{k=-\infty}^\infty e^{-|k|\tau_0} e^{ik\sigma} \, \varphi_n^m(\phi,\sigma) \\
&= D_{nn}^{(m)} \varphi_n^m(\phi,\sigma) +  \frac{P_{n-1/2}^m (z_0) \, Q_{n-1/2}^m (z_0)}{2a} \, \sum_{k=-\infty}^\infty e^{-|k|\tau_0} \, (\cosh\tau_0 - \cos\sigma)^{\frac{3}{2}} \, e^{im\phi} \, e^{i(n+k)\sigma} \\
&= D_{nn}^{(m)} \varphi_n^m(\phi,\sigma) +  \frac{P_{n-1/2}^m (z_0) \, Q_{n-1/2}^m (z_0)}{2a}  \, Q_{n-1/2}^m (z_0) \, \sum_{l=-\infty}^\infty e^{-|l-n|\tau_0} \, (\cosh\tau_0 - \cos\sigma)^{\frac{3}{2}} \, e^{im\phi} \, e^{il\sigma} \\
&= D_{nn}^{(m)} \varphi_n^m(\phi,\sigma) +  \frac{P_{n-1/2}^m (z_0) \, Q_{n-1/2}^m (z_0)}{2a}  \, \sum_{l=-\infty}^\infty e^{-|l-n|\tau_0} \, \frac{\varphi_l^m(\phi,\sigma)}{\Gamma_{ml}} \\
&= D_{nn}^{(m)} \varphi_n^m(\phi,\sigma) + \sum_{l=-\infty}^\infty R_{nl}^{(m)} \varphi_l^m(\phi,\sigma),
\end{align*}
where $R_{nl}^{(m)}$ is given by
\begin{align}
R_{nl}^{(m)}
&= \frac{(-1)^m \sinh\tau_0}{2e^{|l-n|\tau_0}} \frac{\Gamma(l-m+1/2)}{\Gamma(l+m+1/2)} Q_{n-1/2}^m (z_0) P_{l-1/2}^m(z_0) \notag\\
& = \frac{(-1)^m \sinh\tau_0}{2e^{|l-n|\tau_0}}Q_{n-1/2}^m (z_0)  P_{l-1/2}^{-m}(z_0).\label{E}
\end{align}
Here, we used Lemma \ref{toroidrelation} and $z_0 = \cosh\tau_0$.

We now apply Lemma \ref{deriv} to remove derivatives in $D_{nn}^{(m)}$:
\begin{align*}
D_{nn}^{(m)} 
&= \frac{(-1)^m \sinh^2\tau_0}{2} \frac{\Gamma(n-m+1/2)}{\Gamma(n+m+1/2)} \, \Bigg[ Q_{n-1/2}^m(z_0)  \, \frac{\partial}{\partial z} P_{n-1/2}^m (z_0) + P_{n-1/2}^m (z_0) \, \frac{\partial}{\partial z}  Q_{n-1/2}^m (z_0) \Bigg]\\
&= \frac{(-1)^m \sinh^2\tau_0}{2} \Bigg[ Q_{n-1/2}^m(z_0)  \, \frac{\partial}{\partial z} P_{n-1/2}^{-m} (z_0) + P_{n-1/2}^{-m} (z_0) \, \frac{\partial}{\partial z}  Q_{n-1/2}^m (z_0) \Bigg]\\
&= \frac{(-1)^m}{2} \Bigg[ Q_{n-1/2}^m(z_0) \left( (n+m+1/2)P_{n+1/2}^{-m}(z_0) - (n+1/2)  z_0 P_{n-1/2}^{-m}(z_0) \right)  \notag\\
&\qquad\qquad\quad + P_{n-1/2}^{-m} (z_0) \left( (n-m+1/2) Q_{n+1/2}^m(z_0) - (n+1/2)  z_0 Q_{n-1/2}^m(z_0) \right) \Bigg]\notag\\
&= \frac{(-1)^m}{2} \Bigg[ (n-m+1/2) Q_{n+1/2}^m(z_0) P_{n-1/2}^{-m}(z_0) + (n+m+1/2) Q_{n-1/2}^m(z_0) P_{n+1/2}^{-m}(z_0) \notag\\
&\qquad\qquad\quad - (2n+1)  z_0 Q_{n-1/2}^m(z_0) P_{n-1/2}^{-m}(z_0) \Bigg].
\end{align*}
\end{proof}

Let $E = \left[ e^{-|n-l|\tau_0} \right]_{n,l\in\ZZ}$ be a symmetric matrix consisting of exponentials of $\tau_0$ such that
$$
E = 
\begin{bmatrix}
  \ddots & \vdots & \vdots & \vdots & \vdots & \vdots  & \reflectbox{$\ddots$} \\
  \cdots & 1 & e^{-\tau_0} & e^{-2\tau_0} & e^{-3\tau_0} & e^{-4\tau_0}& \cdots \\[3mm]
  \cdots & e^{-\tau_0} & 1 & e^{-\tau_0} & e^{-2\tau_0}& e^{-3\tau_0} & \cdots \\[3mm]
  \cdots & e^{-2\tau_0} & e^{-\tau_0} & 1 & e^{-\tau_0} & e^{-2\tau_0} & \cdots \\[3mm]
  \cdots & e^{-3\tau_0} & e^{-2\tau_0} & e^{-\tau_0} & 1 & e^{-\tau_0} & \cdots \\[3mm]
  \cdots & e^{-4\tau_0} & e^{-3\tau_0} & e^{-2\tau_0} & e^{-\tau_0} & 1 & \cdots \\
  \reflectbox{$\ddots$} & \vdots & \vdots & \vdots & \vdots & \vdots & \ddots
\end{bmatrix}.
$$
For the Kronecker delta function $\delta_{n,l}$,  let $S$ be a shifting matrix, $S = \left[ \delta_{n+1,l} \right]_{n,l\in\ZZ}$, and $\mathcal{Z} = \text{diag}(\ZZ)$ a diagonal matrix such that
$$
S = 
\begin{bmatrix}
  \ddots & \vdots & \vdots & \vdots & \vdots & \vdots  & \reflectbox{$\ddots$} \\
  \cdots & 0& 1 & 0 & 0 & 0 & \cdots \\
  \cdots & 0 & 0 & 1 & 0 & 0 & \cdots \\
  \cdots & 0 & 0 & 0 & 1 & 0 & \cdots \\
  \cdots & 0 & 0 & 0 & 0 & 1 & \cdots \\
  \cdots & 0 & 0 & 0 & 0 & 0 & \cdots \\
  \reflectbox{$\ddots$} & \vdots & \vdots & \vdots & \vdots & \vdots & \ddots
\end{bmatrix},
\qquad 
\mathcal{Z} = 
\begin{bmatrix}
  \ddots & \vdots & \vdots & \vdots & \vdots & \vdots  & \reflectbox{$\ddots$} \\
  \cdots & -2 & 0 & 0 & 0 & 0 & \cdots \\
  \cdots & 0 & -1 & 0 & 0 & 0 & \cdots \\
  \cdots & 0 & 0 & 0 & 0 & 0 & \cdots \\
  \cdots & 0 & 0 & 0 & 1 & 0 & \cdots \\
  \cdots & 0 & 0 & 0 & 0 & 2 & \cdots \\
  \reflectbox{$\ddots$} & \vdots & \vdots & \vdots & \vdots & \vdots & \ddots
\end{bmatrix}.
$$
We also define diagonal matrices, 
$$
\mathcal{Q}^{(m)} = \text{diag}\left(\Big\{Q_{n-1/2}^m(z_0)\Big\}_{n\in\ZZ}\right), 
\qquad \mathcal{P}^{(-m)} = \text{diag}\left(\Big\{P_{n-1/2}^{-m}(z_0)\Big\}_{n\in\ZZ}\right).
$$
In other words, we have
$$
\mathcal{Q}^{(m)}=
\begin{bmatrix}
  \ddots & \vdots & \vdots & \vdots & \vdots & \vdots  & \reflectbox{$\ddots$} \\
  \cdots & Q_{-5/2}^m(z_0) & 0 & 0 & 0 & 0 & \cdots \\[5mm]
  \cdots & 0 & Q_{-3/2}^m(z_0) & 0 & 0 & 0 & \cdots \\[5mm]
  \cdots & 0 & 0& Q_{-1/2}^m(z_0) & 0 & 0 & \cdots \\[5mm]
  \cdots & 0 & 0 & 0 & Q_{1/2}^m(z_0) & 0 & \cdots \\[5mm]
  \cdots & 0 & 0 & 0 & 0 & Q_{3/2}^m(z_0) & \cdots \\
  \reflectbox{$\ddots$} & \vdots & \vdots & \vdots & \vdots & \vdots & \ddots
\end{bmatrix}
$$
and
$$
\mathcal{P}^{(-m)} =
\begin{bmatrix}
  \ddots & \vdots & \vdots & \vdots & \vdots & \vdots  & \reflectbox{$\ddots$} \\
  \cdots & P_{-5/2}^{-m}(z_0) & 0 & 0 & 0 & 0 & \cdots \\[5mm]
  \cdots & 0 & P_{-3/2}^{-m}(z_0) & 0 & 0 & 0 & \cdots \\[5mm]
  \cdots & 0 & 0& P_{-1/2}^{-m}(z_0) & 0 & 0 & \cdots \\[5mm]
  \cdots & 0 & 0 & 0 & P_{1/2}^{-m}(z_0) & 0 & \cdots \\[5mm]
  \cdots & 0 & 0 & 0 & 0 & P_{3/2}^{-m}(z_0) & \cdots \\
  \reflectbox{$\ddots$} & \vdots & \vdots & \vdots & \vdots & \vdots & \ddots
\end{bmatrix}.
$$
The shifting matrix and its transpose give that, for $n\in\ZZ$,
\begin{equation*}
\begin{aligned}
&\left[ S \mathcal{Q}^{(m)} S^t \right]_{n,n} = \left[ \mathcal{Q}^{(m)} \right]_{n+1,n+1} = Q_{n+1/2}^{m}(z_0),\\
&\left[ S \mathcal{P}^{(-m)} S^t \right]_{n,n} = \left[ \mathcal{P}^{(-m)} \right]_{n+1,n+1} = P_{n+1/2}^{-m}(z_0).
\end{aligned}
\end{equation*}
Based on Proposition \ref{Kexpansion} and the above matrix notations, we obtain the matrix representation of the NP operator for tori as the following theorem.
\begin{theorem}[A matrix of the NP operator for tori]\label{NP_matrix}
Let $m\in\ZZ$,  $\tau=\tau_0$, and $z_0 = \cosh\tau_0$.  Let $I$ be the identity matrix, $S$ the shifting matrix, $E$ the expontial matrix, $Z$ the integer diagonal matrix.  Let $\mathcal{Q}^{(m)}$ and $\mathcal{P}^{(-m)}$ be the diagonal matrices related with the associated Legendre functions depending on $z_0$. The matrix representation of $\Kcal_{\p\Om}^*$ along with $\{\varphi_n^m\}_{n\in\ZZ}$, denoted by $\left[\Kcal_{\p\Om}^*\right]^{(m)} $, has the form
\begin{align*}
\left[\Kcal_{\p\Om}^*\right]^{(m)} 
& = \frac{(-1)^m}{2} \Bigg[ \Big(\mathcal{Z}+(1/2-m)I \Big) S\mathcal{Q}^{(m)} S^t \mathcal{P}^{(-m)} + \Big(\mathcal{Z}+(1/2+m)I \Big) \mathcal{Q}^{(m)} S \mathcal{P}^{(-m)} S^t \\
& \qquad\qquad\quad + \sinh\tau_0 \, \mathcal{Q}^{(m)} E \mathcal{P}^{(-m)} - 2\cosh\tau_0 (2Z+I) \mathcal{Q}^{(m)} \mathcal{P}^{(-m)}  \Bigg].
\end{align*}
\end{theorem}

\end{document}